# Einstein from Noise: Statistical Analysis

### v2

@@ I.3 Conditioning on the Fourier frequency noise component
$=\mathbb{E}\left[\left|\mathsf{Y}_{i}[\ell]\right|^{2}\right]\cos\left(\frac{2\pi\ell(r-s)}{d}\right)=\lambda_{\ell}\cos\left(\frac{2\pi\ell(r-s)}{d}\right).$ (I.24)
Substituting (I.24) into (I.22) leads to,
$\mathbb{E}$ $\left[\rho_{k,i}\left[r\right]\rho_{k,i}\left[s\right]|\mathsf{N}_{i}[k]\right]={{\sum_{\ell\in\left\{0,d/2\right\}}\lambda_{\ell}\cdot{\left|\mathsf{X}[\ell]\right|^{2}\,\cos\left(\frac{2\pi\ell}{d}(r-s)\right)}}}$
$+4\cdot{{\sum_{\ell\in\mathcal{I}}\lambda_{\ell}\cdot{\left|\mathsf{X}[\ell]\right|^{2}\,\cos\left(\frac{2\pi\ell}{d}(r-s)\right)}}}.$ (I.25)
@@
${{\sum_{\ell\in\mathcal{I}}{4\left|\mathsf{X}[\ell]\right|^{2}\,\cos\left(\frac{2\pi\ell}{d}(r-s)\right)}}}={{\sum_{\ell\neq\left\{0,k,d/2,d-k\right\}}{2\left|\mathsf{X}[\ell]\right|^{2}\,\cos\left(\frac{2\pi\ell}{d}(r-s)\right)}}}.$ (I.26)
Substituting (I.26) into (I.25), we get,
$\mathbb{E}$ $\left[\rho_{k,i}\left[r\right]\rho_{k,i}\left[s\right]|\mathsf{Y}_{i}[k]\right]=\sum_{\ell=0}^{d-1}\lambda_{\ell}\cdot|\tilde{\mathsf{X}}_{k}[\ell]|^{2}\cos\left(\frac{2\pi\ell}{d}(r-s)\right),$
for $\tilde{\mathsf{X}}_{k}[\ell]$ as defined in (I.14), which completes the proof. ∎
I.4 Convergence of the Fourier phases
Similarly to Appendix B.1 and Lemma B.1, we show here that the imaginary part in (I.8) vanishes. The key observation is that ${\{|\mathsf{Y}_{i}[k]|}\}_{k=0}^{d-1}$, and $\{\phi_{\mathsf{Y}_{i}}[k]\}_{k=0}^{d-1}$ are statistically independent and $\{\phi_{\mathsf{Y}_{i}}[k]\}_{k=0}^{d-1}\sim\mathsf{Unif}[-\pi,\pi)$.
Lemma I.4.
Recall the definition of $\phi_{e,i}[k]$ in (I.5). Then,
$\mu_{\mathsf{A},k}=\mathbb{E}\left[\left|\mathsf{Y}_{1}[k]\right|\sin(\phi_{e,1}[k])\right]=0,$ (I.27)
for every $10\leq k\leq d-$ 1.
Proof of Lemma I.4.
@@ I.4 Convergence of the Fourier phases
$\mathsf{\hat{R}}_{i}=\underset{0\leq r\leq d-1}{\operatorname*{arg\,max}}\ \sum_{k=0}^{d-1}{\left|\mathsf{X}[k]\right|\left|\mathsf{Y}_{i}[k]\right|\,\cos\left(\frac{2\pi kr}{d}+\phi_{\mathsf{Y}_{i}}[k]-\phi_{\mathsf{X}}[k]\right)}.$ (I.28)
@@
$\mathsf{\hat{R}}_{i}=\mathsf{\hat{R}}_{i}\left(\left\{|\mathsf{Y}_{i}[k]|\right\}_{k=0}^{d-1},\left\{|\mathsf{X}[k]|\right\}_{k=0}^{d-1},\left\{\phi_{\mathsf{Y}_{i}}[k]\right\}_{k=0}^{d-1},\left\{\phi_{\mathsf{X}}[k]\right\}_{k=0}^{d-1}\right),$ (I.29)
@@
$\mathsf{\hat{R}}_{1}\left(-\mathsf{D}[0],-\mathsf{D}[1],\ldots,-\mathsf{D}[d-1]\right)=-\mathsf{\hat{R}}_{1}\left(\mathsf{D}[0],\mathsf{D}[1],\ldots,\mathsf{D}[d-1]\right).$ (I.30)
@@
$=\mathbb{E}\left\{\left|\mathsf{Y}_{1}[k]\right|\cdot\left.\mathbb{E}\left[\sin\left(\frac{2\pi k}{d}\mathsf{\hat{R}}_{1}+\phi_{\mathsf{Y}_{1}}[k]-\phi_{\mathsf{X}}[k]\right)\right|\{\left|\mathsf{Y}_{1}[k]\right|\}_{k=0}^{d-1}\right]\right\}.$ (I.31)
The inner expectation in (I.31) is taken w.r.t. the uniform randomness of the phases $\{\phi_{\mathsf{Y}_{1}}[k]\}_{k=0}^{d-1}\in[-\pi,\pi)$. However, due to (I.30), and since the sine function is odd around zero, the integration in (I.31) nullifies. Therefore,
$\left.\mathbb{E}\left[\sin\left(\frac{2\pi k}{d}\mathsf{\hat{R}}_{1}+\phi_{\mathsf{Y}_{1}}[k]-\phi_{\mathsf{X}}[k]\right)\right|\{\left|\mathsf{Y}_{1}[k]\right|\}_{k=0}^{d-1}\right]=0,$ (I.32)
@@
I.5 Convergence to non-vanishing signal
In analogy with Appendix B.2 and Proposition B.2, we now establish that the real part of (I.8) does not vanish.
Proposition I.5.
Recall the definition of $\phi_{e,i}[k]$ in (I.5). Fix $d\in\mathbb{N}$, and assume that $\mathsf{X}[k]\neq 0$ for all $10<k\leq d-$ 1. Then, for any $10\leq k\leq d-$ 1,
$\mu_{\mathsf{B},k}\triangleq\mathbb{E}{\left[\left|\mathsf{Y}_{1}[k]\right|\cos(\phi_{e,1}[k])\right]}>0.$ (I.33)
Proof of Proposition I.5.
@@ I.5 Convergence to non-vanishing signal
$=\mathbb{E}\left[\left.\left|\mathsf{Y}_{1}[k]\right|\cdot\mathbb{E}\left(\cos\left(\frac{2\pi k\mathsf{\hat{R}}_{1}}{d}+\phi_{\mathsf{Y}_{1}}[k]-\phi_{\mathsf{X}}[k]\right)\right|\mathsf{Y}_{1}[k]\right)\right].$ (I.34)
@@
$\frac{1}{2\pi}\int_{0}^{\infty}\mathrm{d}y\ yf_{\left|{\mathsf{Y}_{1}[k]}\right|}(y)\int_{-\pi}^{\pi}\mathrm{d}\varphi\mathbb{E}\left[\left.\cos\left(\frac{2\pi k}{d}\hat{\mathsf{R}}_{1}+\varphi\right)\right|\left|\mathsf{Y}_{1}[k]\right|=y,\phi_{\mathsf{Y}_{1}}[k]=\phi_{\mathsf{X}}[k]+\varphi\right].$ (I.35)
@@
$\qquad\quad+\int_{0}^{\pi}\mathrm{d}\varphi\ \mathbb{E}\left[\left.\cos\left(\frac{2\pi k}{d}\hat{\mathsf{R}}_{1}+\varphi+\pi\right)\right|\left|\mathsf{Y}_{1}[k]\right|=y,\phi_{\mathsf{Y}_{1}}[k]=\phi_{\mathsf{X}}[k]+\varphi+\pi\right].$ (I.36)
@@
$\mathsf{S}^{(+)}=\mathsf{S}_{1}|\mathsf{Y}_{1}[k],$ (I.37)
where the r.h.s. follows from (I.11). By (I.12), the mean vector of $\mathsf{S}_{1}|\mathsf{Y}_{1}[k]$ has a cosine trend, as assumed in Proposition A.3 in (A.34). Additionally, $\mathsf{S}_{1}|\mathsf{Y}_{1}[k]$ is a Gaussian cyclo-stationary process, as described in (I.13). The final condition to verify is that the covariance matrix of $\mathsf{S}_{1}|\mathsf{Y}_{1}[k]$ satisfies $\mathsf{rank}(\mathsf{\Sigma})>d/2$.
By the proposition’s assumption, we assume $\mathsf{X}[k]\neq 0$ for all $10<k\leq d-$ 1, as well as $\mathbb{E}\left[(\mathsf{Y}_{i}[k])^{2}\right]=\lambda_{k}>0$ for every $k\in\left[0,1,\ldots d-1\right]$, which implies $\mathsf{rank}(\mathsf{\Sigma})>d/2$ (see (I.13)). Indeed, for the covariance matrix to have rank greater than $d/2$, at least half of its eigenvalues must be nonzero. As proved in Lemma I.2, the eigenvalues of $\mathsf{\Sigma}$ are given by $\lambda_{\ell}\cdot|\tilde{\mathsf{X}}_{k}[\ell]|^{2}$ for $10\leq\ell\leq d-$ 1. Since the spectrum of $\mathsf{X}$ is non-vanishing, as well as $\lambda_{\ell}>0$, it follows that $\mathsf{rank}(\mathsf{\Sigma})>d/2$. In fact, we obtain $\mathsf{rank}(\mathsf{\Sigma})\geq d-2$, which is larger than $d/2$ for $d>4$. Finally, since the conditional distribution of $\hat{\mathsf{R}}_{1}$ given $\{\left|\mathsf{Y}_{1}[k]\right|=y,\phi_{\mathsf{Y}_{1}}[k]=\phi_{\mathsf{X}}[k]+\varphi\}$ matches that of $\hat{\mathsf{R}}^{(+)}$ in (A.35), and similarly, given $\{\left|\mathsf{Y}_{1}[k]\right|=y,\phi_{\mathsf{Y}_{1}}[k]=\phi_{\mathsf{X}}[k]+\varphi+\pi\}$, it matches $\hat{\mathsf{R}}^{(-)}$ in (A.36), the sum of the integrands on the right-hand side of (I.36) equals the left-hand side of (A.39). By Proposition A.3, this sum is positive for all $\varphi\in[0,\pi]$. Together with (I.35), this completes the proof of Proposition I.5.
I.6 Proof of Proposition 5.4
@@
$\phi_{\mathsf{\hat{X}}}[k]-\phi_{\mathsf{X}}[k]=\arctan\left(\frac{\sum_{i=0}^{M-1}\left|\mathsf{Y}_{i}[k]\right|\sin\left(\phi_{e,i}[k]\right)}{\sum_{i=0}^{M-1}\left|\mathsf{Y}_{i}[k]\right|\cos\left(\phi_{e,i}[k]\right)}\right),$ (I.38)
@@
$\frac{\sum_{i=0}^{M-1}\left|\mathsf{Y}_{i}[k]\right|\sin\left(\phi_{e,i}[k]\right)}{\sum_{i=0}^{M-1}\left|\mathsf{Y}_{i}[k]\right|\cos\left(\phi_{e,i}[k]\right)}\xrightarrow[]{\mathsf{a.s.}}0.$ (I.39)
@@
$\frac{\sum_{i=0}^{M-1}\left|\mathsf{Y}_{i}[k]\right|\sin\left(\phi_{e,i}[k]\right)}{\sum_{i=0}^{M-1}\left|\mathsf{Y}_{i}[k]\right|\cos\left(\phi_{e,i}[k]\right)}\xrightarrow[]{\mathsf{a.s.}}\frac{\mu_{\mathsf{A},k}}{\mu_{\mathsf{B},k}},$ (I.40)
where $\mu_{\mathsf{A},k}\triangleq\mathbb{E}\left[\left|\mathsf{Y}_{1}[k]\right|\sin(\phi_{e,1}[k])\right]$ and $\mu_{\mathsf{B},k}\triangleq\mathbb{E}\left[\left|\mathsf{Y}_{1}[k]\right|\cos(\phi_{e,1}[k])\right]$, defined in (I.9), and (I.10), respectively. By Lemma I.2, $\mu_{\mathsf{A},k}=0$, while by Proposition I.5, we have that $\mu_{\mathsf{B},k}>0$, and thus their ratio converges a.s. to zero by the continuous mapping theorem. Thus, we proved that $\phi_{\mathsf{\hat{X}}}[k]\xrightarrow[]{\mathsf{a.s.}}\phi_{\mathsf{X}}[k]$. Finally, we prove the convergence rate, given in (5.5). According to Proposition B.4, whose assumptions apply for the case of circulant Gaussian noise statistics as well, we have,
$\lim_{M\to\infty}\frac{\mathbb{E}|\phi_{\mathsf{\hat{X}}}[k]-\phi_{\mathsf{X}}[k]|^{2}}{1/M}=\frac{\mathbb{E}\left(\left[\left|\mathsf{Y}_{1}[k]\right|\sin(\phi_{e,1}[k])\right]^{2}\right)}{{\mathbb{E}{\left[\left|\mathsf{Y}_{1}[k]\right|\cos(\phi_{e,1}[k])\right]}}^{2}}<\infty,$ (I.41)

### v3

@@ G.3 Conditioning on the Fourier frequency noise component
$=\frac{1}{2}\mathbb{E}\left[\left|\mathsf{Y}_{i}[\ell]\right|^{2}\right]\cos\left(\frac{2\pi\ell(r-s)}{d}\right)=\frac{\lambda_{\ell}}{2}\cos\left(\frac{2\pi\ell(r-s)}{d}\right).$ (G.27)
Substituting (G.27) into (G.25) leads to,
$\mathbb{E}$ $\left[\rho_{k,i}\left[r\right]\rho_{k,i}\left[s\right]|\mathsf{Y}_{i}[k]\right]={{\sum_{\ell\in\left\{0,d/2\right\}}\frac{\lambda_{\ell}}{2}\cdot{\left|\mathsf{X}[\ell]\right|^{2}\,\cos\left(\frac{2\pi\ell}{d}(r-s)\right)}}}$
$+4\cdot{{\sum_{\ell\in\mathcal{I}}\frac{\lambda_{\ell}}{2}\cdot{\left|\mathsf{X}[\ell]\right|^{2}\,\cos\left(\frac{2\pi\ell}{d}(r-s)\right)}}}.$ (G.28)
@@
$4\cdot{{\sum_{\ell\in\mathcal{I}}{\frac{\lambda_{\ell}}{2}\left|\mathsf{X}[\ell]\right|^{2}\,\cos\left(\frac{2\pi\ell}{d}(r-s)\right)}}}=2{{\sum_{\ell\neq\left\{0,k,d/2,d-k\right\}}{\frac{\lambda_{\ell}}{2}\left|\mathsf{X}[\ell]\right|^{2}\,\cos\left(\frac{2\pi\ell}{d}(r-s)\right)}}}.$ (G.29)
Substituting (G.29) into (G.28), we get,
$\mathbb{E}$ $\left[\rho_{k,i}\left[r\right]\rho_{k,i}\left[s\right]|\mathsf{Y}_{i}[k]\right]=\sum_{\ell=0}^{d-1}\frac{\lambda_{\ell}}{2}\cdot|\widetilde{\mathsf{X}}_{k}[\ell]|^{2}\cos\left(\frac{2\pi\ell}{d}(r-s)\right),$ (G.30)
for $\widetilde{\mathsf{X}}_{k}[\ell]$ as defined in (G.17), which completes the proof. ∎
G.4 Convergence of the Fourier phases
Similarly to Appendix B.1 and Lemma B.1, we show here that the imaginary part in (G.11) vanishes. The key observation is that ${\{|\mathsf{Y}_{i}[k]|}\}_{k=0}^{d-1}$, and $\{\phi_{\mathsf{Y}_{i}}[k]\}_{k=0}^{d-1}$ are statistically independent and $\{\phi_{\mathsf{Y}_{i}}[k]\}_{k=0}^{d-1}\sim\mathsf{Unif}[-\pi,\pi)$.
Lemma G.4.
Recall the definition of $\phi_{e,i}[k]$ in (G.8). Then,
$\mu_{\mathsf{A},k}=\mathbb{E}\left[\left|\mathsf{Y}_{1}[k]\right|\sin(\phi_{e,1}[k])\right]=0,$ (G.31)
for every $0\leq k\leq d-1$.
Proof of Lemma G.4.
@@ G.4 Convergence of the Fourier phases
$\mathsf{\hat{R}}_{i}=\underset{0\leq r\leq d-1}{\operatorname*{arg\,max}}\ \sum_{k=0}^{d-1}{\left|\mathsf{X}[k]\right|\left|\mathsf{Y}_{i}[k]\right|\,\cos\left(\frac{2\pi kr}{d}+\phi_{\mathsf{Y}_{i}}[k]-\phi_{\mathsf{X}}[k]\right)}.$ (G.32)
@@
$\mathsf{\hat{R}}_{i}=\mathsf{\hat{R}}_{i}\left(\left\{|\mathsf{Y}_{i}[k]|\right\}_{k=0}^{d-1},\left\{|\mathsf{X}[k]|\right\}_{k=0}^{d-1},\left\{\phi_{\mathsf{Y}_{i}}[k]\right\}_{k=0}^{d-1},\left\{\phi_{\mathsf{X}}[k]\right\}_{k=0}^{d-1}\right),$ (G.33)
@@
$\mathsf{\hat{R}}_{1}\left(-\mathsf{D}[0],-\mathsf{D}[1],\ldots,-\mathsf{D}[d-1]\right)=-\mathsf{\hat{R}}_{1}\left(\mathsf{D}[0],\mathsf{D}[1],\ldots,\mathsf{D}[d-1]\right).$ (G.34)
@@
$=\mathbb{E}\left\{\left|\mathsf{Y}_{1}[k]\right|\cdot\left.\mathbb{E}\left[\sin\left(\frac{2\pi k}{d}\mathsf{\hat{R}}_{1}+\phi_{\mathsf{Y}_{1}}[k]-\phi_{\mathsf{X}}[k]\right)\right|\{\left|\mathsf{Y}_{1}[k]\right|\}_{k=0}^{d-1}\right]\right\}.$ (G.35)
The inner expectation in (G.35) is taken w.r.t. the uniform randomness of the phases $\{\phi_{\mathsf{Y}_{1}}[k]\}_{k=0}^{d-1}\in[-\pi,\pi)$. However, due to (G.34), and since the sine function is odd around zero, the integration in (G.35) nullifies. Therefore,
$\left.\mathbb{E}\left[\sin\left(\frac{2\pi k}{d}\mathsf{\hat{R}}_{1}+\phi_{\mathsf{Y}_{1}}[k]-\phi_{\mathsf{X}}[k]\right)\right|\{\left|\mathsf{Y}_{1}[k]\right|\}_{k=0}^{d-1}\right]=0,$ (G.36)
@@
G.5 Convergence to non-vanishing signal
In analogy with Appendix B.2 and Proposition B.2, we now establish that the real part of (G.11) does not vanish.
Proposition G.5.
Recall the definition of $\phi_{e,i}[k]$ in (G.8). Fix $d\in\mathbb{N}$, and assume that $\mathsf{X}[k]\neq 0$ for all $0<k\leq d-1$. Then, for any $0\leq k\leq d-1$,
$\mu_{\mathsf{B},k}\triangleq\mathbb{E}{\left[\left|\mathsf{Y}_{1}[k]\right|\cos(\phi_{e,1}[k])\right]}>0.$ (G.37)
Proof of Proposition G.5.
@@ G.5 Convergence to non-vanishing signal
$=\mathbb{E}\left[\left.\left|\mathsf{Y}_{1}[k]\right|\cdot\mathbb{E}\left(\cos\left(\frac{2\pi k\mathsf{\hat{R}}_{1}}{d}+\phi_{\mathsf{Y}_{1}}[k]-\phi_{\mathsf{X}}[k]\right)\right|\mathsf{Y}_{1}[k]\right)\right].$ (G.38)
@@
$\frac{1}{2\pi}\int_{0}^{\infty}\mathrm{d}y\ yf_{\left|{\mathsf{Y}_{1}[k]}\right|}(y)\int_{-\pi}^{\pi}\mathrm{d}\varphi\mathbb{E}\left[\left.\cos\left(\frac{2\pi k}{d}\hat{\mathsf{R}}_{1}+\varphi\right)\right|\left|\mathsf{Y}_{1}[k]\right|=y,\phi_{\mathsf{Y}_{1}}[k]=\phi_{\mathsf{X}}[k]+\varphi\right].$ (G.39)
@@
$\qquad\quad+\int_{0}^{\pi}\mathrm{d}\varphi\ \mathbb{E}\left[\left.\cos\left(\frac{2\pi k}{d}\hat{\mathsf{R}}_{1}+\varphi+\pi\right)\right|\left|\mathsf{Y}_{1}[k]\right|=y,\phi_{\mathsf{Y}_{1}}[k]=\phi_{\mathsf{X}}[k]+\varphi+\pi\right].$ (G.40)
@@
$\mathsf{S}^{(+)}=\mathsf{S}_{1}|\mathsf{Y}_{1}[k],$ (G.41)
where the r.h.s. follows from (G.14). By (G.15), the mean vector of $\mathsf{S}_{1}|\mathsf{Y}_{1}[k]$ has a cosine trend, as assumed in Proposition A.5 in (A.59). Additionally, $\mathsf{S}_{1}|\mathsf{Y}_{1}[k]$ is a Gaussian cyclo-stationary process, as described in (G.16). The final condition to verify is (A.60), which is satisfied by Lemma A.4 (and applies also to circulant Gaussian noise statistics as well).
Since the conditional distribution of $\hat{\mathsf{R}}_{1}$ given $\{\left|\mathsf{Y}_{1}[k]\right|=y,\phi_{\mathsf{Y}_{1}}[k]=\phi_{\mathsf{X}}[k]+\varphi\}$ matches that of $\hat{\mathsf{R}}^{(+)}$ in (A.57), and similarly, given $\{\left|\mathsf{Y}_{1}[k]\right|=y,\phi_{\mathsf{Y}_{1}}[k]=\phi_{\mathsf{X}}[k]+\varphi+\pi\}$, it matches $\hat{\mathsf{R}}^{(-)}$ in (A.58), the sum of the integrands on the right-hand side of (G.40) equals the left-hand side of (A.63). By Proposition A.5, this sum is positive for all $\varphi\in[0,\pi]$. Together with (G.39), this completes the proof of Proposition G.5.
G.6 Proof of Proposition 5.4
@@
$\phi_{\mathsf{\hat{X}}}[k]-\phi_{\mathsf{X}}[k]=\arctan\left(\frac{\sum_{i=0}^{M-1}\left|\mathsf{Y}_{i}[k]\right|\sin\left(\phi_{e,i}[k]\right)}{\sum_{i=0}^{M-1}\left|\mathsf{Y}_{i}[k]\right|\cos\left(\phi_{e,i}[k]\right)}\right),$ (G.42)
@@
$\frac{\sum_{i=0}^{M-1}\left|\mathsf{Y}_{i}[k]\right|\sin\left(\phi_{e,i}[k]\right)}{\sum_{i=0}^{M-1}\left|\mathsf{Y}_{i}[k]\right|\cos\left(\phi_{e,i}[k]\right)}\xrightarrow[]{\mathsf{a.s.}}0.$ (G.43)
@@
$\frac{\sum_{i=0}^{M-1}\left|\mathsf{Y}_{i}[k]\right|\sin\left(\phi_{e,i}[k]\right)}{\sum_{i=0}^{M-1}\left|\mathsf{Y}_{i}[k]\right|\cos\left(\phi_{e,i}[k]\right)}\xrightarrow[]{\mathsf{a.s.}}\frac{\mu_{\mathsf{A},k}}{\mu_{\mathsf{B},k}},$ (G.44)
where $\mu_{\mathsf{A},k}\triangleq\mathbb{E}\left[\left|\mathsf{Y}_{1}[k]\right|\sin(\phi_{e,1}[k])\right]$ and $\mu_{\mathsf{B},k}\triangleq\mathbb{E}\left[\left|\mathsf{Y}_{1}[k]\right|\cos(\phi_{e,1}[k])\right]$, defined in (G.12), and (G.13), respectively. By Lemma G.2, $\mu_{\mathsf{A},k}=0$, while by Proposition G.5, we have that $\mu_{\mathsf{B},k}>0$, and thus their ratio converges a.s. to zero by the continuous mapping theorem. Thus, we proved that $\phi_{\mathsf{\hat{X}}}[k]\xrightarrow[]{\mathsf{a.s.}}\phi_{\mathsf{X}}[k]$. Finally, we prove the convergence rate, given in (5.5). According to Proposition B.4, whose assumptions apply for the case of circulant Gaussian noise statistics as well, we have,
$\lim_{M\to\infty}\frac{\mathbb{E}|\phi_{\mathsf{\hat{X}}}[k]-\phi_{\mathsf{X}}[k]|^{2}}{1/M}=\frac{\mathbb{E}\left(\left[\left|\mathsf{Y}_{1}[k]\right|\sin(\phi_{e,1}[k])\right]^{2}\right)}{{\mathbb{E}{\left[\left|\mathsf{Y}_{1}[k]\right|\cos(\phi_{e,1}[k])\right]}}^{2}}<\infty,$ (G.45)